# On the Fundamental Limits of LLMs at Scale

### v1

@@ 5.2 To What Extent Do LLMs Believe Everything They Read?
In retrieval-augmented LLMs, the same mechanism that enables contextual reasoning, multi-head self-attention introduces a fundamental vulnerability: distraction by irrelevant or misleading context Amiraz et al. (2025); Shi et al. (2023); Yang et al. (2025b). Let the input sequence be a concatenation of the query $q$ and retrieved passages $D_{r}=\{d_{1},\dots,d_{k}\}$. Within each Transformer layer $l$, attention operates as
@@
Empirical analyses confirm this vulnerability: adding a single irrelevant passage can reduce accuracy by up to 30%, with degradation correlating to the distractor’s cosine similarity in embedding space Yang et al. (2025b); Amiraz et al. (2025). Such distractibility arises because pretraining and supervised fine-tuning maximize next-token likelihood conditioned on all tokens’ attention is optimized for coverage, not discrimination. Consequently, attention weights $\alpha_{i}$ do not encode epistemic confidence, and the model’s aggregation of contextual signals remains indiscriminate.
@@
Such training, explored in noise-augmented fine-tuning and entailment-consistency filtering Xiang et al. (2024); Yoran et al. (2023), reduces distraction but at significant computational cost, requiring large curated datasets and layerwise attention attribution. Fundamentally, however, the vulnerability persists: attention aggregates information linearly across all tokens and lacks an intrinsic mechanism to down-weight misinformation. Once a distractor occupies token space within the fixed context window, its representation is fused into the model’s latent state, irreversibly influencing generation. Thus, the same attention mechanism that enables flexible reasoning also guarantees susceptibility to irrelevant or misleading context, a structural limitation of current LLM architectures.
@@
where $\lambda(q)\!\in\![0,1]$ is an implicit trust weight influenced by entity familiarity and prompt framing. Empirically, $\lambda(q)$ varies sharply: models rely on parametric memory for well-known facts but defer to external context when internal confidence is low Du et al. (2024); Wu et al. (2024b).
@@
whose sign determines which source prevails. Because both scores depend on context order and evidence composition, small perturbations in $D_{r}$ or prompt layout can flip $\mathrm{sign}(\Delta)$, producing unpredictable reliance on either source. Benchmarks such as ConflictBank and ClashEval report all four regimes, model-right/context-wrong, model-wrong/context-right, both wrong, and both right across architectures and domains Su et al. (2024); Wu et al. (2024b).
Mechanistically, attention mediates the arbitration: $s_{\text{ctx}}$ emerges from token-level aggregation where retrieved passages accumulate attention mass that can amplify or suppress the parametric signal. Cutting or reweighting early attention heads reduces context dominance when priors are strong, revealing that the LLM’s fusion of sources is coherence-driven, not truth-driven Jin et al. (2024). Even with realistic, time-updated corpora, these behaviors persist, retrieval may override correct priors or reinforce outdated ones Kortukov et al. (2024).
Mitigation strategies fall into three limited families: Training-time biasing involves fine-tuning with conflict-labeled data to calibrate $\lambda(q)$, which is effective but costly and domain-sensitive. Inference-time arbitration uses dual decoding or controller policies to compare parametric and contextual outputs via entailment or consistency, offering robustness yet being latency-heavy and prone to over-deferring to fluent context Wang et al. (2024a). Structure-aware fusion applies graph- or claim-level gating to down-weight $s_{\text{ctx}}$ when contradictions arise, which reduces overt clashes but depends on the quality of structured evidence Kortukov et al. (2024).
@@
causing unstable source arbitration Xu et al. (2024c). Even after calibration or structured fusion, attention still aggregates conflicting signals under likelihood maximization, leaving $\lambda(q)$ ungrounded. Hence, all current LLMs, whether closed-book, open-book, or hybrid, remain vulnerable to contradictions; resolving them requires explicit verifiers or external truth supervision beyond probabilistic conditioning.
@@
A user query $q$ is often an underspecified representation of the underlying intent $z$ Srinivasan et al. (2022). Let $p(z\mid q)$ denote the distribution over possible interpretations and $q(z)$ the disambiguated form. Ideally, retrieval should maximize expected coverage across intents:
@@
but most RAG pipelines collapse this distribution to a single interpretation, approximating $p(z\mid q)\approx\delta(z=z_{0})$. This one-shot formulation biases retrieval toward a single meaning, making the pipeline fragile when $q$ encodes multiple plausible intents. Methods such as RQ-RAG, DMQR-RAG, and Plan-RAG attempt to broaden coverage by generating a set of rewrites $\{q_{i}\}_{i=1}^{m}$, yet rewriting and decomposition introduce new error sources; rewrite quality, selection, and aggregation, and increase computational and ranking complexity Chan et al. (2024); Lee et al. (2024a); Li et al. (2024c).
@@
Hence, ambiguity resolution is implicitly delegated to attention weights rather than an explicit decision over $p(z\mid q)$. If any rewrite retrieves plausible but off-target passages with high $\alpha_{i,j}$, the model’s posterior drifts toward the wrong interpretation even when correct evidence is present. This phenomenon mirrors fragility observed in fusion architectures such as FiD and its variants Izacard and Grave (2020); Ye et al. (2023a).
@@
Empirical studies show that query rewriting, decomposition, and planning improve coverage and attribution on multi-hop and ambiguous queries, yet residual weaknesses persist Chan et al. (2024); Verma et al. (2024). Two structural limitations remain: (1) non-identifiability of intent, because without explicit modeling or user feedback the model cannot determine which latent intent to prioritize, making outcomes sensitive to minor retrieval or prompt changes; and (2) unprincipled fusion, in which attention merges token-level signals without grounding $\alpha_{i,j}$ in a calibrated posterior over intents, thereby optimizing for fluency rather than correctness. Robust handling of query ambiguity, therefore, requires explicit intent inference, decision-aware retrieval, and calibrated context fusion, rather than assuming the LLM’s attention mechanism can self-resolve ambiguity.
@@ 6 When Seeing Fails to Mean Understanding: Limitations of Multimodal Reasoning
By grounding LLMs in perceptual experience, i.e., enabling direct interpretation of visual and auditory inputs, multimodal LLMs (MLLMs) are posited to overcome the hallucinations and abstraction biases that constrain text-only systems (Hurst et al., 2024; Tong et al., 2024a; Wu et al., 2025d). DeepMind’s Flamingo architecture was designed to "bridge powerful pretrained vision-only and language-only models" with the goal of enabling few-shot visual reasoning across diverse tasks without fine-tuning (Alayrac et al., 2022). Similarly, OpenAI’s GPT-4V(ision) was promoted for its "grounded visual understanding" through the ability to "encode, integrate, and reason over arbitrarily interleaved language and vision signals" (Hurst et al., 2024). Google’s Gemini models have positioned multimodality as central to overcoming the grounding problem, i.e., the lack of correspondence between abstract semantic content and real physical objects (Team et al., 2023).
The rationale is intuitive: if text-only models falter from a lack of perceptual grounding, then integrating direct sensory input should alleviate these failures, allowing vision to anchor language and richer modalities to yield more reliable reasoning (Xu et al., 2025c). Yet a paradox emerges from empirical evaluation, i.e., despite richer inputs, MLLMs inherit and often amplify the fundamental limitations of their language-model backbones (Tong et al., 2024a; Cui et al., 2023). The holistic evaluation of GPT-4V shows that regional, language, and prompt framing biases persist, showing that visual input does not eliminate inductive bias from the training data (Cui et al., 2023; Brin et al., 2024; Senkaiahliyan et al., 2023). This is primarily because most MLLMs are built by coupling pretrained vision and language models through modality adapters (Li et al., 2023b; Alayrac et al., 2022). Consequently, the inductive and representational biases embedded in these pretrained components propagate into downstream MLLMs (Tong et al., 2024a), as similarly observed in text models where encoder-level failures transfer to generative tasks (Tong et al., 2023). Furthermore, research on compositional reasoning finds that state-of-the-art MLLMs exhibit the same brittleness documented in text-only models (Ahn et al., 2025; Dziri et al., 2023). The models struggle with systematic generalization, fail on minor perturbations to visual scenes(Zhang et al., 2025e; Clusmann et al., 2025b; a), and collapse when forced to integrate multiple reasoning steps across modalities. Perhaps most troublingly, multimodal architectures introduce novel failure modes, such as visual object hallucinations stemming from "over-reliance on bag-of-objects representations and language priors" (Li et al., 2025c), that are absent in unimodal systems.
@@ 6.1.1 Representation Imbalance: The Hegemony of Text
The feature dimensions between the two components are aligned using a projection layer (Liu et al., 2023b). A two-layer MLP enhancing the vision-language connector representation can improve multimodal capabilities over simple linear projections (Liu et al., 2024b). However, the optimization objective used to train $\mathbf{W_{\text{proj}}}$ is principally text generation, specifically minimizing the causal language modeling loss via modifications in vision to language mapping rather than joint representational adaptation. Empirical analysis reveals the extent of representational imbalance. In VideoLLaMA-7B, output tokens attend to text tokens 157 times more than to visual tokens on a per-token basis (Wu et al., 2025a). This observation of linguistic dominance, driven by dataset and parameter imbalance, can be formalized as the following empirically testable claim
@@ 6.1.4 Tokenization Granularity Mismatch: Quantization and Semantic Distortion
$\mathbf{z}_{q}=\mathbf{e}_{k},\quad k=\arg\min_{j}\|\mathbf{z}_{e}(\mathbf{x})-\mathbf{e}_{j}\|_{2}$
@@ 6.1.5 Advanced Architectures: Mitigations and Persistent Bottlenecks
To address the limitations of simple projection layers, more sophisticated architectures have emerged, though they often rearrange rather than resolve the core issue of linguistic dominance. Query-based bottlenecks like the Q-Former, proposed in BLIP-2 (Li et al., 2023b), use a small set of learnable queries to efficiently extract text-relevant visual features, creating an explicit and highly compressed information channel. Other designs pursue "deep fusion," as seen in DeepStack (Meng et al., 2024), by distributing visual tokens across multiple layers of the LLM. This allows visual information to be processed alongside linguistic information at various depths of the model, enhancing performance, especially on high-resolution tasks. While these approaches enhance performance, the visual information is still processed within a computational structure optimized for text generation. They still operate within a framework that adapts the visual modality to the pre-existing, dominant linguistic structure.
@@ 6.3.1 The Perceptual Illusion of Grounding
MLLMs are built on the hypothesis of perceptual grounding that by processing images (or other modalities) directly, models develop representations anchored in sensory experience rather than purely symbolic abstractions. However, contemporary MLLMs exhibit what we term the perceptual illusion i.e., they appear grounded while continuing to reason over embeddings fundamentally detached from perceptual reality (Cao et al., 2024). The visual embeddings generated are projected into the language model’s token space via $\mathbf{W_{\text{proj}}}$, concatenated with text tokens, and processed by the LLM backbone (1). Critically, the LLM never accesses raw pixel intensities, spatial gradients, or any continuous perceptual features, it operates entirely on discrete token sequences in a learned embedding space (4).
@@
A natural objection is that if MLLMs only generate text, caption-level representations should suffice. However, this conflates output modality with representational requirements. Consider the query: “If this object is rotated 90 degrees clockwise, which face will be visible?” While the answer is text, correct reasoning requires internal representations of 3D geometry and mental rotation, perceptual properties $\mathcal{P}$ underspecified in typical captions $\mathcal{C}$. Formally, let $\mathcal{T}$ denote a task requiring reasoning over $\mathcal{P}$ with textual output. If $I(\mathcal{C};\mathcal{P})<H(\mathcal{P})$, then optimal performance requires $I(\mathcal{R}_{\text{MLLM}};\mathcal{P})>I(\mathcal{C};\mathcal{P})$, contradicting Proposition 5. Caption-mediated representations may memorize common patterns for high in-distribution performance, yet fail on compositional generalization, systematic spatial or causal reasoning, and adversarial robustness. While MLLMs achieve behavioral adequacy (plausible text), they lack representational adequacy (internal perceptual models), evident when tasks demand genuine perceptual reasoning beyond pattern matching. Empirical evidence supports this symbolic detachment, demonstrating that even multimodal models like GPT-4V rely predominantly on textual associations rather than direct visual input when predicting human perceptual judgments (Hirano et al., 2024). The illusion of grounding arises because MLLMs excel at tasks where linguistic priors suffice (e.g., object recognition, scene classification) (Abdou et al., 2021) while failing on tasks requiring genuine perceptual inference (e.g., physical stability, spatial navigation, fine-grained visual reasoning) Rahmanzadehgervi et al. (2025); Hirano et al. (2024); Zhang et al. (2025e); Clusmann et al. (2025a).

### v2

@@ 5.2 To What Extent Do LLMs Believe Everything They Read?
In retrieval-augmented LLMs, the same mechanism that enables contextual reasoning, multi-head self-attention, introduces a fundamental vulnerability: distraction by irrelevant or misleading context (Amiraz et al., 2025; Shi et al., 2023; Yang et al., 2025b). Let the input sequence be a concatenation of the query $q$ and retrieved passages $D_{r}=\{d_{1},\dots,d_{k}\}$. Within each Transformer layer $l$, attention operates as
@@
Empirical analyses confirm this vulnerability: adding a single irrelevant passage can reduce accuracy by up to 30%, with degradation correlating to the distractor’s cosine similarity in embedding space (Yang et al., 2025b; Amiraz et al., 2025). Such distractibility arises because pretraining and supervised fine-tuning maximize next-token likelihood conditioned on all tokens’ attention is optimized for coverage, not discrimination. Consequently, attention weights $\alpha_{i}$ do not encode epistemic confidence, and the model’s aggregation of contextual signals remains indiscriminate.
@@
Such training, explored in noise-augmented fine-tuning and entailment-consistency filtering (Xiang et al., 2024; Yoran et al., 2023), reduces distraction but at significant computational cost, requiring large curated datasets and layerwise attention attribution. Fundamentally, however, the vulnerability persists: attention aggregates information linearly across all tokens and lacks an intrinsic mechanism to down-weight misinformation. Once a distractor occupies token space within the fixed context window, its representation is fused into the model’s latent state, irreversibly influencing generation. Thus, the same attention mechanism that enables flexible reasoning also guarantees susceptibility to irrelevant or misleading context, a structural limitation of current LLM architectures.
@@
where $\lambda(q)\!\in\![0,1]$ is an implicit trust weight influenced by entity familiarity and prompt framing. Empirically, $\lambda(q)$ varies sharply: models rely on parametric memory for well-known facts but defer to external context when internal confidence is low (Du et al., 2024; Wu et al., 2024b).
@@
whose sign determines which source prevails. Because both scores depend on context order and evidence composition, small perturbations in $D_{r}$ or prompt layout can flip $\mathrm{sign}(\Delta)$, producing unpredictable reliance on either source. Benchmarks such as ConflictBank and ClashEval report all four regimes, model-right/context-wrong, model-wrong/context-right, both wrong, and both right across architectures and domains (Su et al., 2024; Wu et al., 2024b).
Mechanistically, attention mediates the arbitration: $s_{\text{ctx}}$ emerges from token-level aggregation where retrieved passages accumulate attention mass that can amplify or suppress the parametric signal. Cutting or reweighting early attention heads reduces context dominance when priors are strong, revealing that the LLM’s fusion of sources is coherence-driven, not truth-driven (Jin et al., 2024). Even with realistic, time-updated corpora, these behaviors persist, retrieval may override correct priors or reinforce outdated ones (Kortukov et al., 2024).
Mitigation strategies fall into three limited families: Training-time biasing involves fine-tuning with conflict-labeled data to calibrate $\lambda(q)$, which is effective but costly and domain-sensitive. Inference-time arbitration uses dual decoding or controller policies to compare parametric and contextual outputs via entailment or consistency, offering robustness yet being latency-heavy and prone to over-deferring to fluent context (Wang et al., 2024a). Structure-aware fusion applies graph- or claim-level gating to down-weight $s_{\text{ctx}}$ when contradictions arise, which reduces overt clashes but depends on the quality of structured evidence (Kortukov et al., 2024).
@@
causing unstable source arbitration (Xu et al., 2024c). Even after calibration or structured fusion, attention still aggregates conflicting signals under likelihood maximization, leaving $\lambda(q)$ ungrounded. Hence, all current LLMs, whether closed-book, open-book, or hybrid, remain vulnerable to contradictions; resolving them requires explicit verifiers or external truth supervision beyond probabilistic conditioning.
@@
A user query $q$ is often an underspecified representation of the underlying intent $z$ (Srinivasan et al., 2022). Let $p(z\mid q)$ denote the distribution over possible interpretations and $q(z)$ the disambiguated form. Ideally, retrieval should maximize expected coverage across intents:
@@
but most RAG pipelines collapse this distribution to a single interpretation, approximating $p(z\mid q)\approx\delta(z=z_{0})$. This one-shot formulation biases retrieval toward a single meaning, making the pipeline fragile when $q$ encodes multiple plausible intents. Methods such as RQ-RAG, DMQR-RAG, and Plan-RAG attempt to broaden coverage by generating a set of rewrites $\{q_{i}\}_{i=1}^{m}$, yet rewriting and decomposition introduce new error sources; rewrite quality, selection, and aggregation, and increase computational and ranking complexity (Chan et al., 2024; Lee et al., 2024a; Li et al., 2024c).
@@
Hence, ambiguity resolution is implicitly delegated to attention weights rather than an explicit decision over $p(z\mid q)$. If any rewrite retrieves plausible but off-target passages with high $\alpha_{i,j}$, the model’s posterior drifts toward the wrong interpretation even when correct evidence is present. This phenomenon mirrors fragility observed in fusion architectures such as FiD and its variants (Izacard and Grave, 2020; Ye et al., 2023a).
@@
Empirical studies show that query rewriting, decomposition, and planning improve coverage and attribution on multi-hop and ambiguous queries, yet residual weaknesses persist (Chan et al., 2024; Verma et al., 2024). Two structural limitations remain: (1) non-identifiability of intent, because without explicit modeling or user feedback the model cannot determine which latent intent to prioritize, making outcomes sensitive to minor retrieval or prompt changes; and (2) unprincipled fusion, in which attention merges token-level signals without grounding $\alpha_{i,j}$ in a calibrated posterior over intents, thereby optimizing for fluency rather than correctness. Robust handling of query ambiguity, therefore, requires explicit intent inference, decision-aware retrieval, and calibrated context fusion, rather than assuming the LLM’s attention mechanism can self-resolve ambiguity.
@@ 6 When Seeing Fails to Mean Understanding: Limitations of Multimodal Reasoning
By grounding LLMs in perceptual experience, i.e., enabling direct interpretation of visual and auditory inputs, multimodal LLMs (MLLMs) are posited to overcome the hallucinations and abstraction biases that constrain text-only systems (Hurst et al., 2024; Tong et al., 2024a; Wu et al., 2025d). DeepMind’s Flamingo architecture was designed to “bridge powerful pretrained vision-only and language-only models” with the goal of enabling few-shot visual reasoning across diverse tasks without fine-tuning (Alayrac et al., 2022). Similarly, OpenAI’s GPT-4V(ision) was promoted for its “grounded visual understanding” through the ability to “encode, integrate, and reason over arbitrarily interleaved language and vision signals” (Hurst et al., 2024). Google’s Gemini models have positioned multimodality as central to overcoming the grounding problem, i.e., the lack of correspondence between abstract semantic content and real physical objects (Team et al., 2023).
The rationale is intuitive: if text-only models falter from a lack of perceptual grounding, then integrating direct sensory input should alleviate these failures, allowing vision to anchor language and richer modalities to yield more reliable reasoning (Xu et al., 2025c). Yet a paradox emerges from empirical evaluation, i.e., despite richer inputs, MLLMs inherit and often amplify the fundamental limitations of their language-model backbones (Tong et al., 2024a; Cui et al., 2023). The holistic evaluation of GPT-4V shows that regional, language, and prompt framing biases persist, showing that visual input does not eliminate inductive bias from the training data (Cui et al., 2023; Brin et al., 2024; Senkaiahliyan et al., 2023). This is primarily because most MLLMs are built by coupling pretrained vision and language models through modality adapters (Li et al., 2023b; Alayrac et al., 2022). Consequently, the inductive and representational biases embedded in these pretrained components propagate into downstream MLLMs (Tong et al., 2024a), as similarly observed in text models where encoder-level failures transfer to generative tasks (Tong et al., 2023). Furthermore, research on compositional reasoning finds that state-of-the-art MLLMs exhibit the same brittleness documented in text-only models (Ahn et al., 2025; Dziri et al., 2023). The models struggle with systematic generalization, fail on minor perturbations to visual scenes(Zhang et al., 2025e; Clusmann et al., 2025b; a), and collapse when forced to integrate multiple reasoning steps across modalities. Perhaps most troublingly, multimodal architectures introduce novel failure modes, such as visual object hallucinations stemming from “over-reliance on bag-of-objects representations and language priors” (Li et al., 2025b), that are absent in unimodal systems.
@@ 6.1.1 Representation Imbalance: The Hegemony of Text
The feature dimensions between the two components are aligned using a projection layer (Liu et al., 2023b). A two-layer MLP enhancing the vision-language connector representation can improve multimodal capabilities over simple linear projections (Liu et al., 2024b). However, the optimization objective used to train $\mathbf{W_{\text{proj}}}$ is principally text generation, specifically minimizing the causal language modeling loss via modifications in vision to language mapping rather than joint representational adaptation. Empirical analysis reveals the extent of representational imbalance. In VideoLLaMA-7B, output tokens attend to text tokens 157 times more than to visual tokens on a per-token basis (Wu et al., 2025a). This observation of linguistic dominance, driven by dataset and parameter imbalance, can be formalized as the following empirically testable claim.
@@ 6.1.4 Tokenization Granularity Mismatch: Quantization and Semantic Distortion
$\mathbf{z}_{q}=\mathbf{e}_{k},\quad k=\arg\min_{j}\|\mathbf{z}_{e}(\mathbf{x})-\mathbf{e}_{j}\|_{2}.$
@@ 6.1.5 Advanced Architectures: Mitigations and Persistent Bottlenecks
To address the limitations of simple projection layers, more sophisticated architectures have emerged, though they often rearrange rather than resolve the core issue of linguistic dominance. Query-based bottlenecks like the Q-Former, proposed in BLIP-2 (Li et al., 2023b), use a small set of learnable queries to efficiently extract text-relevant visual features, creating an explicit and highly compressed information channel. Other designs pursue “deep fusion,” as seen in DeepStack (Meng et al., 2024), by distributing visual tokens across multiple layers of the LLM. This allows visual information to be processed alongside linguistic information at various depths of the model, enhancing performance, especially on high-resolution tasks. While these approaches enhance performance, the visual information is still processed within a computational structure optimized for text generation. They still operate within a framework that adapts the visual modality to the pre-existing, dominant linguistic structure.
@@ 6.3.1 The Perceptual Illusion of Grounding
MLLMs are built on the hypothesis of perceptual grounding that by processing images (or other modalities) directly, models develop representations anchored in sensory experience rather than purely symbolic abstractions. However, contemporary MLLMs exhibit what we term the perceptual illusion, i.e., they appear grounded while continuing to reason over embeddings fundamentally detached from perceptual reality (Cao et al., 2024). The visual embeddings generated are projected into the language model’s token space via $\mathbf{W_{\text{proj}}}$, concatenated with text tokens, and processed by the LLM backbone (1). Critically, the LLM never accesses raw pixel intensities, spatial gradients, or any continuous perceptual features; it operates entirely on discrete token sequences in a learned embedding space (4).
@@
A natural objection is that if MLLMs only generate text, caption-level representations should suffice. However, this conflates output modality with representational requirements. Consider the query: “If this object is rotated 90 degrees clockwise, which face will be visible?” While the answer is text, correct reasoning requires internal representations of 3D geometry and mental rotation, perceptual properties $\mathcal{P}$ underspecified in typical captions $\mathcal{C}$. Formally, let $\mathcal{T}$ denote a task requiring reasoning over $\mathcal{P}$ with textual output. If $I(\mathcal{C};\mathcal{P})<H(\mathcal{P})$, then optimal performance requires $I(\mathcal{R}_{\text{MLLM}};\mathcal{P})>I(\mathcal{C};\mathcal{P})$, contradicting Proposition 5. Caption-mediated representations may memorize common patterns for high in-distribution performance, yet fail on compositional generalization, systematic spatial or causal reasoning, and adversarial robustness. While MLLMs achieve behavioral adequacy (plausible text), they lack representational adequacy (internal perceptual models), evident when tasks demand genuine perceptual reasoning beyond pattern matching. Empirical evidence supports this symbolic detachment, demonstrating that even multimodal models like GPT-4V rely predominantly on textual associations rather than direct visual input when predicting human perceptual judgments (Hirano et al., 2024). The illusion of grounding arises because MLLMs excel at tasks where linguistic priors suffice (e.g., object recognition, scene classification) (Abdou et al., 2021) while failing on tasks requiring genuine perceptual inference (e.g., physical stability, spatial navigation, fine-grained visual reasoning) (Rahmanzadehgervi et al., 2025; Hirano et al., 2024; Zhang et al., 2025e; Clusmann et al., 2025a).